# Vanilla Bayesian Optimization Performs Great in High Dimensions

### 25 Feb 2024 v3

@@ 1 Introduction
    We prove that when the model is uninformed,  EI will not exihibit exploratory behavior along the boundary, contrasting claims of (Swersky, 2017; Oh et al., 2018)
@@ 4.3 Complexity of Existing HDBO
Having established that high complexity can yield uninformative models, as well as having discussed the link between complexity and dimensionality, it is evident that complexity assumptions must be sensible to facilitate a calibrated GP in HDBO. Notably, all the classes of HDBO algorithms outlined in Sec. 3 have such complexity-lowering assumptions. In Fig. 4, we display the modeled complexity of the most common classes of HDBO algorithms. Fig. 4 can be viewed as a cross section of Fig. 2, where we fix the number of data points to 1000 and instead vary the dimensionality of the problem to demonstrate how each HDBO class lower the growth of the complexity in the dimensionality, relative to a common global GP model as well as an independent kernel. The methods presented are: REMBO (Wang et al., 2016) with $d_{e}=4$, random AddGPs (Ziomek & Bou Ammar, 2023), BOCK (Oh et al., 2018), Local GPs (Eriksson et al., 2019) after one round of shrinkage, the global GP with fixed lengthscales from Fig. 2, and our proposed method — scaling the lengthscales in the dimensionality of the problem, which is introduced later in Sec. 5. While each algorithm has parameters that affect its MIG, we have set parameters to make the comparison as fair as possible. In App. A.2, we outline the setup in detail.
@@ 5 Low-complexity High-dimensional
Hypothesizing that the pitfalls of HDBO are strictly caused by assumptions of insurmountable complexity, we present our main methodological contribution. We design simple, plug-and-play assumption that retains almost constant complexity as the dimensionality increases. Similar to Fig. 1, we achieve this by adjusting the prior on the lengthscales to the dimensionality of the problem to the task at hand. Moreover, we ensure a calibrated signal variance by drawing on previous findings on GPs in an over-parameterized regime.
@@ 5.2 Calibrating Epistemic Uncertainty
which, in a BO context, has a different impact than in GPs generally. As we are able to selectively acquire our data, a large number of parameters and substantially correlated data will simplify data fit, driving down the optimal value of $\sigma_{f}^{2}$. When data is repeatedly re-normalized, the issue will be reinforced, as the signal variance is further decreased, and another highly correlated query is selected. As such, we fix $\sigma_{f}^{2}=1$ to match the scale of the standardized observations, and to ensure that $\sigma_{f}^{2}$ does not diminish over time. In App. D.4, we ablate this modeling choice and demonstrate its impact on performance.
@@ 6 Results
We now compare vanilla BO, equipped with the complexity-reducing assumptions from Sec. 5, against state-of-the-art methods across various classes of HDBO: the subspace-methods BAxUS, Bounce and SAASBO (Papenmeier et al., 2022, 2023; Eriksson & Jankowiak, 2021), the Local BO algorthms TuRBO (Eriksson et al., 2019) and Maximal Probability of Descent (Nguyen et al., 2022) (MPD), the AddGP method RD-UCB (Ziomek & Bou Ammar, 2023), the variable selection method MCTS-VS (Song et al., 2022), and CMA-ES (Hansen, 2006). We use each method’s official repository, with the exception of SAASBO which is run through Ax (Bakshy et al., ). We detail the experimental setup in App. A, and our code is publically available at https://github.com/hvarfner/vanilla_bo_in_highdim.
We instantiate Vanilla BO with $\mu_{0}=\sqrt{2},\sigma_{0}=\sqrt{3}$, which equates to $\bm{\ell}\approx 0.50$ for $D=6$. We initialize our method with $3\sqrt{D}$ SOBOL samples to acquire substantial data for modeling. We initialize the baselines with 30 samples, marked by the dashed vertical line. We note that some methods (BAxUS, CMA-ES, MPD) deviate from conventional initialization schemes. On all benchmarks, we run Vanilla BO with LogEI (Ament et al., 2023), using a low acquisition optimization budget of 512 initial (global) SOBOL samples and 512 Gaussian samples around the incumbent, followed by L-BFGS on the 4 best candidates, which is made possible by the low-complexity-high-smoothness model. In App. D, we demonstrate that our method delivers consistent performance for a range of complexities.
@@ 6.1 Sparse Synthetic Test Functions
We start by evaluating Vanilla BO on a collection of commonly considered synthetic test functions with varying total and effective dimensionality. We note that the assumptions made in Sec. 5 diametrically oppose these test cases - each function has a low number of highly important dimensions with the large remainder being unimportant, whereas we assume that each dimension has relatively small impact.
Vanilla BO is highly performant, as it rapidly identifies the important dimensions and subsequently optimizes the task. This is similar in to (Eriksson & Jankowiak, 2021), whose assumptions, represented through its sparse lengthscale prior, aligns perfectly with the task at hand. As such, SAASBO should, and does, perform best on average, with Vanilla BO being second. Notably, as Vanilla BO does not require the HMC (Bingham et al., 2018) fully Bayesian model fitting that SAASBO uses, it runs in a small fraction of the time.
@@ 6.2 High-dimensional Optimization Tasks
We now benchmark Vanilla BO against a collection of frequently considered tasks in the high-dimensional literature (Eriksson & Jankowiak, 2021; Eriksson et al., 2019; Papenmeier et al., 2022, 2023; Šehić et al., 2021): Specifically, we consider MOPTA08 (124D), SVM (388D), Lasso-DNA (180D), and the MuJoCo (Todorov et al., 2012) Ant ( [^8] D) and Humanoid (6392D) reinforcement learning tasks. We stress that, for all benchmarks where applicable, (BAxUS on SVM, Lasso-DNA and Humanoid, TuRBO on MOPTA and SVM, RD-UCB on Lasso-DNA), baselines perform within the error bars of the original implementation (Papenmeier et al., 2022; Ziomek & Bou Ammar, 2023) or in other papers by the same authors in the case of TuRBO (Eriksson et al., 2019; Eriksson & Jankowiak, 2021).
Vanilla BO is extremely competitive, performing the best by a substantial margin on Lasso-DNA, SVM and Humanoid. On Ant, BAxUS’ low-dimensional initialization allows it to obtain an average value of 450 after DoE due to consistently sampling data points close to the center of the search space. On MOPTA, Vanilla BO is outperformed only by SAASBO, and outperforms the other methods by a similarly large amount. Notably, Vanilla BO is very consistent between repetitions, as evident by the small error bars. This can be attributed to the consistent modelling, as Vanilla BO is not dependent on randomness in subspace design, trust region initialization or variable selection. Rather, it obtains a consistent, calibrated model through reasonable correlation in the data, from which it can effectively infer promising regions and quickly improve upon the DoE. The effective inference capabilities are most apparent on SVM, nt and Humanoid, where Vanilla BO rapidly improves after DoE (60, 90 and 240 iterations respectively). In App. D.1, we compare to conventional MAP and MLE, and demonstrate that the potent inference ability is unique to our method.
@@ 7 Conclusion and Future Work
Nevertheless, we do not believe that tailored high-dimensional BO algorithms are unwarranted: if the problem at hand is known to adhere to the structural assumptions that are conventionally made (effective subspace, additivity) or where non-stationarity in the objective facilitates local modelling, we believe these approaches will be superior to the vanilla algorithm. However, these restrictive assumptions should not be made solely out of necessity, but rather when prior knowledge supports them. For future work, we plan to investigate the topic of complexity as it relates to modelling in Bayesian optimization more broadly. Specifically, we believe latent space GP models (Griffiths & Hernández-Lobato, 2017; Maus et al., 2022) can incorporate complexity considerations to enable more efficient modeling.
Broader Impact
@@ C.1 Theory: Non-existence of the Boundary Issue
To prove that EI will never prefer a point with maximal variance under the proposed setup, we first restate the formal definition of EI: For a noiseless function, the next point $\bm{x}_{*}$, where $y_{max}$ is the maximal objective function value observed so far, as
$\bm{x}_{*}\in\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}\mathbb{E}\left[[(f(\bm{x})-y_{\text{max}}]^{+}\right]=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}Zs_{n}(\bm{x})\Phi(Z)+s_{n}(\bm{x})\phi(Z),$ (6)
where $Z=(\mu(\bm{x})-y_{\text{max}})/\sigma(\bm{x})$, and $\phi$ and $\Phi$ are the PDF and CDF of the standard Gaussian, respectively. Note that we are considering a maximization problem. Lastly, we denote the mean of the GP by $c$.
@@
[…]
@@
Moreover, for ease of notation, we scale our data such that $c=0$ and $\sigma_{f}^{2}=1$, $\hat{y}_{\text{max}}=\frac{y_{\text{max}}-c}{\sigma_{f}}>0$, $Z=\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right)$ and
$\texttt{EI}{}(\rho)=\left(\rho-1\right)\hat{y}_{\text{max}}\Phi\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right)+\sqrt{1-\rho^{2}}\phi\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right).$ (9)
@@
$\frac{\partial\texttt{EI}{}}{\partial\rho}=\hat{y}_{\text{max}}\Phi(Z)-\frac{\rho}{\sqrt{1-\rho^{2}}}\phi(Z).$ (10)
We then establish the values of $\rho$ for which $\frac{\partial\texttt{EI}{}}{\partial\rho}$ are positive. We bound $\Phi(Z)$ as a function of $\phi(Z)$ using known inequalities for Mills’ ratio $(1-\Phi)/\phi$ (Duembgen, 2010),
$\frac{2\phi(Z)}{\sqrt{4+Z^{2}}-Z}\leq\Phi(Z),\quad-\hat{y}_{\text{max}}\leq Z\leq 0.$
We then insert the inequality into the expression for $\frac{\partial\texttt{EI}{}}{\partial\rho}$, which leaves
$\frac{\partial\texttt{EI}{}}{\partial\rho}$ $\geq\left[\frac{2\hat{y}_{\text{max}}}{\sqrt{4+Z^{2}}-Z}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(Z)$
$=\left[\frac{2\hat{y}_{\text{max}}}{\sqrt{4+\frac{\left(1-\rho\right)^{2}}{1-\rho^{2}}\hat{y}_{\text{max}}^{2}}+\frac{1-\rho}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(Z)$
$\geq\left[\frac{2\hat{y}_{\text{max}}}{{2+2\hat{y}_{\text{max}}\frac{1-\rho}{\sqrt{1-\rho^{2}}}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(Z),$
where we apply the triangle inequality to the denominator of the left term between the first and second step. We wish to lower bound $\rho^{*}=\mathop{\mathrm{arg\,max}}\texttt{EI}(\rho)$ by finding the values of $\rho$ for which EI is strictly increasing. Thus, we set our lower bound for the derivative to zero and obtain
$\frac{\partial\texttt{EI}{}}{\partial\rho}\geq\left[\frac{\hat{y}_{\text{max}}}{{1+\hat{y}_{\text{max}}\frac{1-\rho}{\sqrt{1-\rho^{2}}}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(Z)\geq 0.$
We then multiply by $\sqrt{1-\rho^{2}}$, and divide by $\phi(Z)$, which are both $>0$ for $\rho>0$, and re-arrange:
@@
As such, we can conclude, that EI is strictly increasing in $\rho$ as long as ${\hat{y}_{\text{max}}}\geq\rho{\sqrt{\frac{1+\rho}{1-\rho}}}$, meaning that $\rho^{*}=\mathop{\mathrm{arg\,max}}_{\rho\in[0,1)}\texttt{EI}({\rho})$ satisfies the inequality

### 26 Jun 2024 v4

@@ 1 Introduction
    We prove that when the model is uninformed,  EI will not exihibit exploratory behavior along the boundary, contrasting claims of (Swersky, 2017; Oh et al., 2018).
@@ 4.3 Complexity of Existing HDBO
Having established that high complexity can yield uninformative models, as well as having discussed the link between complexity and dimensionality, it is evident that complexity assumptions must be sensible to facilitate a calibrated GP in HDBO. Notably, all the classes of HDBO algorithms outlined in Sec. 3 have such complexity-lowering assumptions. In Fig. 4, we display the modeled complexity of the most common classes of HDBO algorithms. Fig. 4 can be viewed as a cross section of Fig. 2, where we fix the number of data points to 1000 and instead vary the dimensionality of the problem to demonstrate how each HDBO class lower the growth of the complexity in the dimensionality, relative to a common global GP model as well as an independent kernel. The methods presented are: REMBO (Wang et al., 2016) with $d_{e}=4$, random AddGPs (Ziomek & Bou Ammar, 2023), BOCK (Oh et al., 2018), Local GPs (Eriksson et al., 2019) after one round of shrinkage, the global GP with fixed lengthscales from Fig. 2, and our proposed method — scaling the lengthscales in the dimensionality of the problem, which is introduced later in Sec. 5. While each algorithm has parameters that affect the MIG, we have set parameters to make the comparison as fair as possible. In App. A.2, we outline the setup in detail.
@@ 5 Low-complexity High-dimensional
Hypothesizing that the pitfalls of HDBO are strictly caused by assumptions of insurmountable complexity, we present our main methodological contribution. We design a simple, plug-and-play assumption that retains almost constant complexity as the dimensionality increases. Similar to Fig. 1, we achieve this by adjusting the prior on the lengthscales to the dimensionality of the problem to the task at hand. Moreover, we ensure a calibrated signal variance by drawing on previous findings on GPs in an over-parameterized regime.
@@ 5.2 Calibrating Epistemic Uncertainty
which, in a BO context, has a different impact than in GPs generally. As we are able to selectively acquire our data, a large number of parameters and substantially correlated data will simplify data fit, driving down the optimal value of $\sigma_{f}^{2}$. When data is repeatedly re-normalized, the issue will be reinforced, as the signal variance is further decreased, and another highly correlated query is selected. As such, we fix $\sigma_{f}^{2}=1$ to match the scale of the standardized observations, and to ensure that $\sigma_{f}^{2}$ does not diminish over time. In App. F.1, we ablate this modeling choice and demonstrate its impact on the value of the outputscale during optimization for two real-world optimization tasks, and in App. F.2, we demonstrate the impact of the outputscale modeling choice on optimization performance.
@@ 6 Results
We now compare our Vanilla Bayesian Optimization method with a dimensionality-scaled lengthscale prior (we will refer to our method as $D$-scaled $p(\bm{\ell)}$ or DSP for clarity), against state-of-the-art HDBO methods. We will include various classes of HDBO methods, such as the subspace-methods Bounce and SAASBO (Papenmeier et al., 2023; Eriksson & Jankowiak, 2021), the Local BO algorithms TuRBO (Eriksson et al., 2019) and Maximal Probability of Descent (Nguyen et al., 2022) (MPD), the AddGP method RD-UCB (Ziomek & Bou Ammar, 2023), the variable selection method MCTS-VS (Song et al., 2022), and CMA-ES (Hansen, 2006). We use each method’s official repository, with the exception of SAASBO which is run through Ax (Bakshy et al., ). We detail the experimental setup in App. A, and our code is publicly available at https://github.com/hvarfner/vanilla_bo_in_highdim.
We instantiate the DSP with $\mu_{0}=\sqrt{2},\sigma_{0}=\sqrt{3}$, which equates to $\bm{\ell}\approx 0.50$ for $D=6$ under the mode of $p(\bm{\ell})$. We initialize all methods with 30 samples, marked by a dashed vertical line. Bounce, CMA-ES and MPD deviate from conventional initialization. On all benchmarks, we use LogEI (Ament et al., 2023), using a low acquisition optimization budget of 512 initial (global) SOBOL samples and 512 Gaussian samples around the incumbent, followed by L-BFGS on the 4 best candidates, which is made possible by the low-complexity-high-smoothness model. In App. E, we demonstrate that our method delivers consistent performance for a range of complexities.
Figure 6: Best observed value for the DSP, conventional MAP, and TuRBO on Hartmann (6D) and four mid-dimensional real-world optimization tasks. All methods perform comparably on Hartmann, while the DSP outperforms or is on part with TuRBO on the other tasks. Notably, the DSP performs at least equally well as the $\Gamma(3,6)$ on all tasks, and substantially better on 4 out of 5 tasks, which suggests that it is well-suited as a drop-in replacement for conventional priors.
@@ 6.1 Sparse Synthetic Test Functions
We start by evaluating the DSP on a collection of commonly considered synthetic test functions with varying total and effective dimensionality. We note that the assumptions made in Sec. 5 diametrically oppose these test cases - each function has a low number of highly important dimensions with the large remainder being unimportant, whereas we assume that each dimension has relatively small impact.
The DSP is highly performant, as it rapidly identifies the important dimensions and subsequently optimizes the task. This is similar in to (Eriksson & Jankowiak, 2021), whose assumptions, represented through its sparse lengthscale prior, aligns perfectly with the task at hand. As such, SAASBO should, and does, perform best on average, with Vanilla BO being second. Notably, as Vanilla BO does not require the HMC (Bingham et al., 2018) fully Bayesian model fitting that SAASBO uses, it runs in a small fraction of the time.
6.2 A Plug-in on Mid-Dimensional Tasks
Subsequently, we use the DSP as a plug-in for low- and mid-dimensional tasks, primarily those considered in (Eriksson et al., 2019), to evaluate its ability to serve as a substitute for conventional, non-adaptive hyperparameter priors. The Lunar Lander (12D) and Robot Pushing (14D) tasks from (Wang et al., 2017), as well as the Swimmer (16D) and Hopper (32D) reinforcement learning tasks from the MuJoCo suite (Todorov et al., 2012), where we aim to learn a linear policy for two objects with varying degrees of freedom. We evaluate against a $\Gamma(3,6)$ lengthscale prior with learnable $\sigma^{2}_{f}$, and against TuRBO, commonly considered the state-of-the-art mid-dimensional BO method. In Fig. 6, it is shown that the DSP is either competitive with, or outperforms, TuRBO on all tasks. In App. C, we empirically display the pronounced local search-like behavior of the $\Gamma(3,6)$ prior for Hartmann (6D), which suggests that it does not offer a calibrated exploration-exploitation trade-off. On the contrary, the DSP maintains a moderate distance between queries, which indicates a calibrated trade-off throughout optimization.
@@ 6.3 High-dimensional Optimization Tasks
We now benchmark Vanilla BO with the DSP against a collection of frequently considered tasks in the high-dimensional literature (Eriksson & Jankowiak, 2021; Eriksson et al., 2019; Papenmeier et al., 2022, 2023; Šehić et al., 2021): Specifically, we consider MOPTA08 (124D), SVM (388D), Lasso-DNA (180D), and the MuJoCo (Todorov et al., 2012) Ant ( [^8] D) and Humanoid (6392D) reinforcement learning tasks. We stress that, for all benchmarks where applicable, (BAxUS on SVM, Lasso-DNA and Humanoid, TuRBO on MOPTA and SVM, RD-UCB on Lasso-DNA), baselines perform within the error bars of the original implementation (Papenmeier et al., 2022; Ziomek & Bou Ammar, 2023) or in other papers by the same authors in the case of TuRBO (Eriksson et al., 2019; Eriksson & Jankowiak, 2021). In App. D, we display the performance in a lower-budget setting. In App. E.1, we compare to conventional MAP and MLE.
Figure 8: Distribution of lengthscale values for the DSP on the 388D SVM task. Lengthscales are sorted according to their mean log value. The three last indexed dimensions (385, 386, 387) are considered particularly active (Eriksson & Jankowiak, 2021), but are not identified as such consistently in our method. the black horizontal lines indicate upper and lower quartiles, and the orange horizontal lines indicate the median.
Fig. 7 shows that Vanilla BO with the DSP is highly competitive, performing the best by a substantial margin on Lasso-DNA and Humanoid, and produces top-two performance on the remaining tasks. On the MuJuCo Ant, Bounce’s low-dimensional initialization allows it to obtain an average value of 800 after DoE due to consistently sampling data points close to the center of the search space. Notably, the DSP is very consistent between repetitions, as evident by the small error bars. This can be attributed to the consistent modelling, as the DSP is not dependent on randomness in subspace design, trust region initialization or variable selection. Rather, it obtains a consistent, calibrated model through meaningful correlation in the data, from which it can effectively infer promising regions and improve upon the DoE.
Notably, the DSP does not heavily depend on identification of active variables. In Fig. 8, we demonstrate for the distribution of lengthscale values for the 388D SVM after 250 iterations. The DSP does not consistently identify active dimensions with large confidence. Instead, the calibrated complexity of the model allows for meaningful inference along all dimensions, until particularly active dimensions are potentially identified. As such, the DSP does not require the identification of active variables to achieve calibrated BO, but its identification helps optimization. As such, we attribute the superior performance of our method to the calibrated complexity, and the effective inference and exploration-exploitation trade-off that stems from it.
@@ 7 Conclusion and Future Work
Nevertheless, we do not believe that tailored high-dimensional BO algorithms are unwarranted: if the problem at hand is known to adhere to the structural assumptions that are conventionally made (effective subspace, additivity) or where non-stationarity in the objective facilitates local modelling, we believe these approaches will be superior to the vanilla algorithm. However, these restrictive assumptions should not be made out of necessity, but when prior knowledge supports them. For future work, we plan to investigate the topic of complexity as it relates to modelling in Bayesian optimization more broadly, and in the context of latent space GP models (Griffiths & Hernández-Lobato, 2017; Maus et al., 2022).
Impact Statement
@@
Acknowledgements
Carl Hvarfner, Erik Hellsten and Luigi Nardi were partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Luigi Nardi was partially supported by the Wallenberg Launch Pad (WALP) grant nbr. 2021.0348.
@@ C.1 Theory: Non-existence of the Boundary Issue
To prove that EI will never prefer a point with maximal variance under the proposed setup, we first restate the formal definition of EI:
$\texttt{EI}{}(\bm{x})=\mathbb{E}\left[[(f(\bm{x})-y_{\text{max}}]^{+}\right]=Z\sigma(\bm{x})\Phi(Z)+\sigma(\bm{x})\phi(Z),$ (6)
where $y_{max}$ is the maximal objective function value observed so far, $Z=(\mu(\bm{x})-y_{\text{max}})/\sigma(\bm{x})$, and $\phi$ and $\Phi$ are the PDF and CDF of the standard Gaussian, respectively. Hence, for a noiseless function the next point $\bm{x}_{*}$ to query is
$\bm{x}_{*}\in\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}\texttt{EI}{}(\bm{x}).$ (7)
Note that we are considering a maximization problem. Lastly, we denote the mean of the GP by $c$.
@@
Assume that $y_{\text{max}}>c$, $\mathbf{K}=\bm{I}$ and that the candidate query $\bm{x}_{*}$ correlates with at most one observation. Then, the correlation $\rho^{*}=\sigma_{f}^{-2}k(\bm{x}_{*},\bm{x}_{\text{inc}})$ between the next query $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}\texttt{EI}{}(\bm{x})$ and $\bm{x}_{\text{inc}}$ satisfies
$\rho^{*}\sqrt{\frac{1+\rho^{*}}{1-\rho^{*}}}\geq\frac{y_{\text{max}}-c}{\sigma_{f}}.$ (8)
@@
We will by $\hat{\texttt{EI}{}}(\rho)$ denote the $\rho$-parametrized EI, which is a function over $\rho$, to distinguish it from the regular $\texttt{EI}{}(\bm{x})$, which is a function over $\bm{x}$. Moreover, for ease of notation, we scale our data such that $c=0$, $\sigma_{f}^{2}=1$, and $\hat{y}_{\text{max}}=\frac{y_{\text{max}}-c}{\sigma_{f}}>0$. After scaling we further get $\hat{Z}=\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right)$, and finally:
$\hat{\texttt{EI}{}}(\rho)=\left(\rho-1\right)\hat{y}_{\text{max}}\Phi\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right)+\sqrt{1-\rho^{2}}\phi\left(\frac{\left(\rho-1\right)}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}\right).$ (10)
@@
$\frac{\partial\hat{\texttt{EI}{}}}{\partial\rho}=\hat{y}_{\text{max}}\Phi(\hat{Z})-\frac{\rho}{\sqrt{1-\rho^{2}}}\phi(\hat{Z}).$ (11)
We then establish the values of $\rho$ for which $\frac{\partial\texttt{EI}{}}{\partial\rho}$ are positive. We bound $\Phi(\hat{Z})$ as a function of $\phi(\hat{Z})$ using known inequalities for Mills’ ratio $(1-\Phi)/\phi$ (Duembgen, 2010),
$\frac{2\phi(\hat{Z})}{\sqrt{4+\hat{Z}^{2}}-\hat{Z}}\leq\Phi(\hat{Z}),\quad-\hat{y}_{\text{max}}\leq\hat{Z}\leq 0.$
We then insert the inequality into the expression for $\frac{\partial\hat{\texttt{EI}{}}}{\partial\rho}$, which leaves
$\frac{\partial\hat{\texttt{EI}{}}}{\partial\rho}$ $\geq\left[\frac{2\hat{y}_{\text{max}}}{\sqrt{4+\hat{Z}^{2}}-\hat{Z}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(\hat{Z})$
$=\left[\frac{2\hat{y}_{\text{max}}}{\sqrt{4+\frac{\left(1-\rho\right)^{2}}{1-\rho^{2}}\hat{y}_{\text{max}}^{2}}+\frac{1-\rho}{\sqrt{1-\rho^{2}}}\hat{y}_{\text{max}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(\hat{Z})$
$\geq\left[\frac{2\hat{y}_{\text{max}}}{{2+2\hat{y}_{\text{max}}\frac{1-\rho}{\sqrt{1-\rho^{2}}}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(\hat{Z}),$
where we apply the triangle inequality to the denominator of the left term between the first and second step. We wish to lower bound $\rho^{*}=\mathop{\mathrm{arg\,max}}\hat{\texttt{EI}}(\rho)$ by finding the values of $\rho$ for which $\hat{\texttt{EI}}$ is strictly increasing. Thus, we set our lower bound for the derivative to zero and obtain
$\frac{\partial\hat{\texttt{EI}{}}}{\partial\rho}\geq\left[\frac{\hat{y}_{\text{max}}}{{1+\hat{y}_{\text{max}}\frac{1-\rho}{\sqrt{1-\rho^{2}}}}}-\frac{\rho}{\sqrt{1-\rho^{2}}}\right]\phi(\hat{Z})\geq 0.$
We then multiply by $\sqrt{1-\rho^{2}}$, and divide by $\phi(\hat{Z})$, which are both $>0$ for $\rho>0$, and re-arrange:
@@
As such, we can conclude, that $\hat{\texttt{EI}{}}$ is strictly increasing in $\rho$ as long as ${\hat{y}_{\text{max}}}\geq\rho{\sqrt{\frac{1+\rho}{1-\rho}}}$, meaning that $\rho^{*}=\mathop{\mathrm{arg\,max}}_{\rho\in[0,1)}\hat{\texttt{EI}}({\rho})$ satisfies the inequality